# INTERVENOR[Uncaptioned image]: Prompt the Coding Ability of Large Language Models with the Interactive Chain of Repairing

### 19 Jan 2024 v3

@@ Abstract
This paper proposes INTERactiVE chaiN Of Repairing (INTERVENOR), which mimics human code repairing behavior (iteratively judging, rethinking, and repairing) and prompts the coding ability of regard Large Language Models (LLMs). Specifically, INTERVENOR employs two LLM based agents, Code Learner and Code Teacher, to play different roles in code repairing and work interactively to repair the generated codes. The Code Learner is asked to generate and repair code according to the instructions from the Code Teacher. The Code Teacher rethinks the code errors according to the corresponding feedback from compilers and iteratively generates the chain-of-repairing (CoR) to guide the code repairing process for Code Learner. Our experiments show that INTERVENOR outperforms baseline models and achieves about 13% and 4.5% improvements over GPT-3.5 in code generation and code translation tasks, respectively. Our further analyses show that CoR can illuminate the bug reasons and solution plans via natural language. With the feedback of code compilers, INTERVENOR can accurately identify the syntax errors and assertion errors and provide precise instructions to repair codes. All data and codes are available at https://github.com/NEUIR/INTERVENOR.
INTERVENOR[Uncaptioned image]: Prompt the Coding Ability of Large Language Models with the Interactive Chain of Repairing
@@ 1 Introduction
Large Language Models (LLMs), such as ChatGPT OpenAI (2022), have demonstrated extremely strong emergency ability and benefited lots of tasks, especially for code generation OpenAI (2023); Roziere et al. (2023); Wang et al. (2023). It greatly stimulates the efficiency and productivity of coding and software development Qian et al. (2023a). Existing code based models focus on pretraining language models on code corpus Muennighoff et al. (2023); Luo et al. (2023); Li et al. (2023b); Zheng et al. (2023) and using Chain-of-Thought (CoT) to prompt the coding ability of LLMs Wei et al. (2023); Huang et al. (2023); Li et al. (2023a). Nevertheless, forcing LLMs to directly generate the totally correct coding results is challenging even for skilled programmers in real-world scenarios Chen et al. (2023).
@@
Figure 1: The Illustration of INTERVENOR. There are two agents in INTERVENOR, the teacher and student, whom collaborate to fix the code. The error messages are utilized as a kind of INTERVENORRefer to caption to alleviate the Degeneration-of-Thought (DoT) problem.
@@
Recently, researchers focus on self-improving code generation methods using Self-Repair methods Olausson et al. (2023); Chen et al. (2023), which employ LLMs to execute codes and repair codes by themselves for improving the quality of generated codes. Moreover, multi-agent collaborative coding approaches Qian et al. (2023a); Dong et al. (2023) have also proven their effectiveness in handling difficult code tasks by prompting LLMs to play different roles, such as developers and testers. However, lots of bugs are difficult to find due to the cognitive inertia McGuire (1960)–overlooking the buggy codes that may not conform to their pre-existing coding thinking of LLMs. Thus, these agent based code refine methods Dong et al. (2023); Qian et al. (2023a) are highly dependent on the self-evaluation ability of LLMs and may face the Degeneration-of-Thought (DoT) problem Liang et al. (2023); Shinn et al. (2023).
This paper proposes INTERactiVE chaiN Of Repairing (INTERVENOR) to intervene the DoT problem in code repairing by incorporating the feedback from code compilers to facility the bug detection. However, there is a significant gap between compiler errors and task instructions Wang et al. (2022b) for code repair. Thus, INTERVENOR proposes an interactive Chain-of-Repairing (CoR) mechanism to bridge the gap from compiler feedback and repair instruction. Follow Dong et al. (2023), we design two agents Code Learner and Code Teacher to repair codes. As shown in Figure 1, the Code Learner is responsible for both initial code generation and iterative code repair, while the Code Teacher iteratively generates the CoR to guide the code repairing process for Code Learner. Different from Self-Debug Chen et al. (2023), Code Teacher leverages the bug report from compilers and the program generated by Code Learner to rethink and generate insightful error explanations and comprehensive planning on how to address the errors. This interactive code repair process will continue until the code learner successfully fixes all the code errors or reaches the predetermined maximum number of repair attempts. Besides, we collect the buggy codes that are generated from GPT-3.5 and build the CodeError dataset to facilitate the code repairing related tasks.
@@
Our experiments demonstrate the effectiveness of INTERVENOR by outperforming the state-of-the-art code generation and code translation models. Notably, INTERVENOR also achieves about 13% and 4.5% improvements over the GPT-3.5 model OpenAI (2022) in code generation and code translation tasks, showing its ability to improve the quality of generated codes via iterative code repairing. Our further analyses illustrate that INTERVENOR is effective in leveraging the bug messages from code compilers, recognizing the reasons for code errors, and also providing correction planning in natural language. Thanks to our CoR mechanism, INTERVENOR avoids thinking by LLMs themselves and can accurately diagnose the buggy codes and correct the assertion errors and name/syntax errors even in more difficult code generation scenarios. Our CoR mechanism enables LLMs to avoid designing complex prompts for generating codes and achieve the best performance via only three-turn code repairing. INTERVENOR shows the potential in leveraging the feedback from environments or rule systems to evolve LLMs Olausson et al. (2023).
@@ 2 Related Work
The code generation tasks Chen et al. (2021); Austin et al. (2021); Zheng et al. (2023) aim to generate correct and executable code based on the given natural language description, which has drawn lots of attention from researchers. LLMs such as ChatGPT OpenAI (2022) and GPT-4 OpenAI (2023) have shown strong effectiveness in generating code of high quality. To enhance the coding ability of LLMs, existing work focuses on code-specific pretraining and performs exceptionally well in code generation tasks Roziere et al. (2023); Li et al. (2023b); Luo et al. (2023); Wang et al. (2023).
@@
Some models focus on utilizing some prompting techniques to enhance the coding capabilities of LLMs. CodeCoT Huang et al. (2023) is inspired by Chain-of-Thought Wei et al. (2023) and prompts the quality of generated codes using Code-CoT and the Self-exam CodeCoT methods. LLMs are asked to craft the code and design a set of test cases to generate and polish the codes. Structured Chain-of-Thought (SCoT) Li et al. (2023a) further considers the program structure, such as sequences, branches, and loops, and prompts LLMs to generate intermediate reasoning steps with program structures. Nevertheless, forcing LLMs to directly generate the correct program results is challenging in the code generation task Chen et al. (2023).
@@
To better polish the generated codes, existing work focuses on Self-Refine Olausson et al. (2023) and Self-Repair Chen et al. (2023) methods. The Self-Refine models improve the quality of generated codes LLMs by decoding multiple samples and then selecting the most suitable one based on certain criteria. To build tailored criteria, one approach is to execute the generated codes and select the most optimal one based on the execution results Ni et al. (2023); Zhang et al. (2022); Shi et al. (2022); Li et al. (2022). Another approach is to rerank multiple code solutions to select the final code Shi et al. (2022); Zhang et al. (2022); Chen et al. (2022); Inala et al. (2022). However, these methods need to consume a lot of computing resources to generate code candidates, which is inefficient Zhang et al. (2023).
@@
The other research direction is utilizing an iterative code repair approach to further improve the quality of generated codes Zhang et al. (2023); Welleck et al. (2022); Madaan et al. (2023); Shinn et al. (2023). Self-Debug Chen et al. (2023) utilizes the explanations generated by LLMs to fix self-generated code, whereas Self-Repair Olausson et al. (2023) employs human-provided feedback for improvement. Self-Edit Zhang et al. (2023) uses error messages to refine the generated code, but it requires training an additional fault-aware editor, which involves additional complex tasks such as model design and construction of training data. Instead of directly incorporating feedback for code repair, INTERVENOR designs an additional agent to rethink the reasons for coder errors and generate the chain-of-repairing to teach the other agent to repair codes via natural language.
@@
Moreover, the work Dong et al. (2023); Qian et al. (2023a) also designs a multi-agent collaborative approach to simulate the software development process and improve the efficiency of code generation. Nevertheless, these methods are highly dependent on the self-evaluation ability of LLMs and may face the Degeneration-of-Thought (DoT) problem Liang et al. (2023); Shinn et al. (2023). Unlike them, INTERVENOR focuses on the bug-fixing process and proposes a simple but effective solution, which utilizes external tools Qian et al. (2023b), such as the Python interpreter, to execute the code Xu et al. (2023); Qian et al. (2023b) and use the accurate bug report to facility the agent collaboration during the interactive code repairing.
@@
Figure 2: Illustration of Our Interactive Chain of Repairing Model (INTERVENOR Refer to caption).
@@ 3 Methodology
In this section, we introduce INTERVENOR, which conducts an interactive program repairing using LLM collaboration. We first describe the preliminary of code repairing (Sec. 3.1) and then introduce our interactive Chain-of-Repairing (CoR) mechanism (Sec. 3.2).
3.1 Preliminary of Code Repairing
The code repairing models mainly focus on self-repair Olausson et al. (2023); Chen et al. (2023). These models usually consist of three steps, including code generation, code execution, and code explanation. Self-repair aims to use LLM itself to conduct the self-debug and self-execution processes and then iteratively repair codes. Nevertheless, programmers usually fail to recognize the code errors because of cognitive inertia, making the code execution more difficult by LLM itself.
Different from these self-repair models, INTERVENOR follows previous work Zhang et al. (2023); Wang et al. (2022a) to incorporate the feedback from compilers to prompt the code generation ability of LLMs. Instead of directly feeding the code bug message to LLMs, we design the interactive chain of repairing mechanism, which builds two agents to rethink and repair the code errors. The compiler serves as an INTERVENOR to avoid the Degeneration-of-Thought (DoT) problem during the interactive repairing process.
3.2 Interactive Chain-of-Repairing (CoR)
As shown in Figure 2, given the coding task, INTERVENOR aims to mimic the human bug repairing behavior by iteratively acquiring feedback from compilers, then generating bug reasons and solving plans in natural language, and finally fixing the program. Specifically, INTERVENOR employs two LLM based agents to play different roles in code repairing (Sec. 3.2.1). Then it conducts an interactive code repairing process with the agents (Sec. 3.2.2).
@@ 3.2.1 Agent Building
INTERVENOR involves the integration of two agents, Code Learner and Code Teacher, who work interactively to repair the generated codes. The Code Learner generates codes and repairs codes according to the instruction from Code Teacher. The Code Teacher focuses more on rethinking and illuminating the code errors for students. More agent building details are shown in Appendix A.4.
Code Learner. The code learner conducts two coding behaviors, including first time code generation and interactive code repairing.
In the first time code generation process, the code learner aims to produce the initial code version according to the requirements of the given coding task. Then we change the agent role for code repairing and use the instruction “You are a student assistant with excellent code repair capabilities” to activate the code repairing ability of LLMs. Such an instruction aims to stimulate the ability of LLMs to conduct execution by themselves, rethink the code errors according to the code repair instruction from the code teacher, and fix the bugs.
Code Teacher. The code teacher plays a critical role in interactive code repairing. It aims to generate the code error explanation and bug fix planning for the code learner to help it repair codes.
Specifically, we use the prompt “You are an experienced and insightful programming instructor, and you need to identify the bugs in the given code based on the error messages.” to activate the bug tracing ability and code diagnosis ability. Subsequently, the teacher will offer detailed repair suggestions and guidance to help the student better comprehend and resolve the errors in the code.
3.2.2 Interactive Code Repairing Workflow
INTERVENOR conducts an interactive code repairing process and we will introduce the communication among agents and code compiler.
@@ 4 Experimental Methodology
In this section, we describe the datasets, evaluation metrics, baselines, and implementation details in our experiments.
@@
Table 1: Statistics of the Data Used in Our Experiments. #Tests represents the average number of test cases.
@@
Dataset. We evaluate the code generation and translation effectiveness on three datasets, including HumanEval, MBPP and HumanEval-X. Besides, we build a new benchmark CodeError to further test the code repairing ability. All data statistics are shown in Table 1.
@@
HumanEval Chen et al. (2021) serves as a benchmark for evaluating the functional correctness of synthesized programs generated from docstrings. It comprises 164 hand-written Python programming problems, which consist of function signatures, docstrings, bodies, and multiple unit tests. MBPP Austin et al. (2021) is a benchmark that includes 974 introductory-level Python programming problems. Each problem comprises a problem statement, a code solution, and three automated test cases, and the task IDs that range from 11 to 510 are used for evaluation. HumanEval-X  Zheng et al. (2023) is a benchmark that is used to assess a model’s multi-programming language generation and translation capability. It consists of 820 human-crafted data instances, covering C++, Java, JavaScript, and Go.
Then we build the CodeError benchmark to further evaluate the code repairing effectiveness of INTERVENOR. We collect data from HumanEval and MBPP and reserve the instances, which conduct code errors in the code generation experiment. Specifically, for these code generation tasks, we utilize GPT-3.5 to generate codes according to the task instructions. These generated codes usually contain various bugs and the code errors include syntax errors, logical errors, and others. In this paper, we use the Python interpreter to test the generated code and reserve the buggy codes that do not pass the Python interpreter. These buggy codes and error messages are used to construct the CodeError benchmark. We provide more details about the CodeError benchmark in Appendix  A.2.
@@
Evaluation Metrics. We use $Pass@k$ Chen et al. (2021) to evaluate the code generation and translation effectiveness of different models, which is the same as the previous work Chen et al. (2021); Zheng et al. (2023); Li et al. (2023a); Chen et al. (2022); Nijkamp et al. (2023).
@@
Baselines. We first compare INTERVENOR with several code-oriented large language pretrained models, such as Incoder Fried et al. (2022), CodeGen Nijkamp et al. (2023), CodeGeeX Zheng et al. (2023), CodeT5 Wang et al. (2023), StarCoder Li et al. (2023b), WizardCoder Luo et al. (2023), and Llama based models Touvron et al. (2023); Roziere et al. (2023). These models are pretrained on large-scale code corpora, demonstrating strong code generation capabilities. Additionally, we also compare INTERVENOR with some closed-source and high-performance large language models, e.g. Claude Anthropic (2023), GPT-3.5 OpenAI (2022), and GPT-4 OpenAI (2023), which show strong emergent abilities, especially for the code generation tasks. In our experiments, GPT-3.5 is our main baseline model. Besides, we also compare Self-Debug Chen et al. (2023) and the multi-agent collaborative method, Self-Collaboration Dong et al. (2023) to show the code repairing ability of INTERVENOR.
@@
Implementation Details. In our experiments, we use GPT-3.5 as the foundation model to build different agents in our INTERVENOR model. We set the temperature to 0.2 and the maximum generation length to 512 tokens. The maximum number of interactive code repairs is set to 5.

### 20 Feb 2024 v4

# INTERVENOR[Uncaptioned image]: Prompting the Coding Ability of Large Language Models with the Interactive Chain of Repair

@@ Abstract
This paper introduces INTERVENOR (INTERactiVE chaiN Of Repair), a system designed to emulate the interactive code repair processes observed in humans, encompassing both code diagnosis and code repair. INTERVENOR prompts Large Language Models (LLMs) to play distinct roles during the code repair process, functioning as both a Code Learner and a Code Teacher. Specifically, the Code Learner is tasked with adhering to instructions to generate or repair code, while the Code Teacher is responsible for crafting a Chain-of-Repair (CoR) to serve as guidance for the Code Learner. During generating the CoR, the Code Learner needs to check the generated codes from Code Learner and reassess how to address code bugs based on error feedback received from compilers. Experimental results demonstrate that INTERVENOR surpasses baseline models, exhibiting improvements of approximately 18% and 4.3% over GPT-3.5 in code generation and code translation tasks, respectively. Our further analyses show that CoR is effective to illuminate the reasons behind bugs and outline solution plans in natural language. With the feedback of code compilers, INTERVENOR can accurately identify syntax errors and assertion errors and provide precise instructions to repair codes. All data and codes are available at https://github.com/NEUIR/INTERVENOR.
INTERVENOR[Uncaptioned image]: Prompting the Coding Ability of Large Language Models with the Interactive Chain of Repair
@@ 1 Introduction
Large Language Models (LLMs), such as ChatGPT OpenAI (2022), have shown remarkable performance on code related tasks OpenAI (2023); Roziere et al. (2023); Wang et al. (2023b). This has significantly enhanced the efficiency and productivity in coding and software development Qian et al. (2023a). Current approaches for code-based models involve pretraining language models on code corpora Muennighoff et al. (2023); Luo et al. (2023); Li et al. (2023b); Zheng et al. (2023) and employing Chain-of-Thought (CoT) to prompt the coding proficiency of LLMs Wei et al. (2022); Huang et al. (2023); Li et al. (2023a). However, compelling LLMs to directly generate entirely correct code proves to be exceptionally challenging, even for proficient programmers in real-world scenarios Chen et al. (2023b).
@@
Figure 1: The Illustration of INTERVENOR. There are two agents in INTERVENOR, the teacher and student, who collaborate to repair the code. The error messages are utilized as a kind of INTERVENORRefer to caption to alleviate the Degeneration-of-Thought (DoT) problem.
@@
Recently, researchers focus on improving the code generation ability of LLMs through Self-Repair techniques Olausson et al. (2023); Chen et al. (2023b). These methods leverage LLMs themselves to execute and repair codes, thereby enhancing the quality of generated code. Moreover, multi-agent collaborative coding approaches Qian et al. (2023a); Dong et al. (2023) have also proven their effectiveness in handling difficult code tasks by prompting LLMs to play different roles, such as developers and testers. However, lots of bugs are difficult to find due to the cognitive inertia McGuire (1960)–overlooking the buggy codes that may not conform to their pre-existing coding thinking of LLMs. These agent-based code refinement methods Dong et al. (2023); Qian et al. (2023a) heavily rely on the self-evaluation capabilities of LLMs, potentially encountering the Degeneration-of-Thought (DoT) problem Liang et al. (2023); Shinn et al. (2023).
This paper proposes INTERactiVE chaiN Of Repair (INTERVENOR) to alleviate the DoT problem in code repair. The approach incorporates feedback from code compilers to enhance the code repair process. Following Dong et al. (2023), we develop two agents, namely, Code Learner and Code Teacher, to collaboratively repair codes interactively. As illustrated in Figure 1, the Code Learner is tasked with generating/repairing codes based on provided instructions. To craft specific code repair instructions to guide the Code Learner Kaddour et al. (2023); Wang et al. (2023a), the Code Teacher generates the Chain-of-Repair (CoR) to illustrate the bug repair solutions for the Code Learner. Instead of Self-Debug Chen et al. (2023b), Code Teacher incorporates the bug report from compilers to rethink the reasons of code errors and generate planning on how to repair the bugs. This interactive code repair process will continue until the code learner successfully fixes all code errors or reaches the predetermined maximum number of repair attempts.
Experimental results demonstrate the effectiveness of INTERVENOR by outperforming previous baseline models. Notably, INTERVENOR also achieves about 18% and 4.3% improvements over GPT-3.5 OpenAI (2022) in code generation and code translation tasks, showing its ability to improve the quality of generated codes through iterative code repair. Besides, we also build the CodeError dataset for evaluating the code repair ability of INTERVENOR by collecting the buggy code snippets from GPT-3.5 and real-world user-submitted codes. INTERVENOR further validates its efficacy by demonstrating a twofold increase in the number of successfully repaired codes.
@@
Our further analyses illustrate that INTERVENOR is effective in leveraging the bug messages from code compilers, recognizing the reasons for code errors, and providing correction planning in natural language. Thanks to our CoR mechanism, INTERVENOR avoids thinking by LLMs themselves and can accurately diagnose the buggy codes and correct the assertion errors and name/syntax errors even in more difficult code generation scenarios. Our CoR mechanism enables LLMs to avoid designing complex code generation/repair prompts and achieve the best performance via only three-turn code repair. It also shows the potential to leverage the feedback from environments or rule systems to evolve LLMs Olausson et al. (2023).
@@ 2 Related Work
Code generation tasks Chen et al. (2021); Austin et al. (2021); Zheng et al. (2023) aim to generate correct and executable code based on the given natural language description, which has drawn lots of attention from researchers. LLMs such as ChatGPT OpenAI (2022) and GPT-4 OpenAI (2023) have shown strong effectiveness in generating code of high quality. To enhance the coding ability of LLMs, existing work focuses on code-specific pretraining and performs exceptionally well in code generation tasks Roziere et al. (2023); Li et al. (2023b); Luo et al. (2023); Wang et al. (2023b).
@@
Recently, some models focus on utilizing prompting techniques to enhance the coding capabilities of LLMs. CodeCoT Huang et al. (2023) is inspired by Chain-of-Thought Wei et al. (2022) and prompts the quality of generated codes using Code-CoT and Self-exam methods. LLMs are asked to craft the code and design a set of test cases to polish the codes. Structured Chain-of-Thought (SCoT) Li et al. (2023a) further considers the program structure, such as sequences, branches, and loops, and prompts LLMs to generate intermediate reasoning steps with program structures. Nevertheless, forcing LLMs to directly generate the completely correct codes is challenging in the code generation task Chen et al. (2023b).
@@
To generate more accurate code, existing efforts primarily concentrate on Self-Refine Olausson et al. (2023) and Self-Repair Chen et al. (2023b) techniques. The Self-Refine models aspire to improve the quality of generated code by decoding multiple samples and subsequently selecting the most suitable one based on specific criteria. One strategy to formulate customized criteria involves executing the generated code and selecting the optimal one based on the resulting execution outcomes Ni et al. (2023); Zhang et al. (2023b); Shi et al. (2022); Li et al. (2022). Another approach is to rerank multiple code solutions to determine the final code Shi et al. (2022); Zhang et al. (2023b); Chen et al. (2023a); Inala et al. (2022). However, these methods necessitate significant computing resources for generating code candidates, rendering them inefficient Zhang et al. (2023a).
@@
Another research avenue involves employing an iterative code repair approach to enhance the quality of generated code Zhang et al. (2023a); Welleck et al. (2023); Madaan et al. (2023); Shinn et al. (2023). Self-Debug Chen et al. (2023b) utilizes explanations generated by LLMs to rectify self-generated code, while Self-Repair Olausson et al. (2023) incorporates human-provided feedback for improvement. Self-Edit Zhang et al. (2023a) employs error messages to refine generated code, but it necessitates the training of an additional fault-aware editor to generate a new program. Instead of directly integrating feedback for code repair, INTERVENOR designs an additional agent to reflect the reasons for coder errors and generate the Chain-of-Repair (CoR). This CoR is then employed to instruct the other agent in code repair through natural language.
@@
Moreover, the work Dong et al. (2023); Qian et al. (2023a) also designs a multi-agent collaborative approach to simulate the software development process and improve the efficiency of code generation. Nevertheless, these methods are highly dependent on the self-evaluation ability of LLMs and may face the Degeneration-of-Thought (DoT) problem Liang et al. (2023); Shinn et al. (2023). Unlike them, INTERVENOR focuses on the bug-fixing process and proposes a simple but effective solution, which utilizes external tools, such as the Python interpreter, to execute the code Xu et al. (2023); Qian et al. (2023b) and use the accurate bug report to facilitate the agent collaboration during interactive code repair.
@@
Figure 2: Illustration of Our Interactive Chain-of-Repair Model (INTERVENOR Refer to caption).
@@ 3 Methodology
In this section, we introduce INTERVENOR, which conducts an interactive program repair process using LLM collaboration. We first describe the preliminary of code repair (Sec. 3.1) and then introduce our interactive Chain-of-Repair (CoR) mechanism (Sec. 3.2).
3.1 Preliminary of Code Repair
The code repair models mainly focus on Self-Repair Olausson et al. (2023); Chen et al. (2023b). These models usually consist of three steps, including code generation, code execution, and code explanation. Self-Repair aims to use LLM itself to conduct the self-debug and self-execution processes and then iteratively repair codes. Nevertheless, programmers usually fail to recognize the code errors because of cognitive inertia McGuire (1960), making the code execution more difficult by LLM itself.
Different from these self-repair models, INTERVENOR follows previous work Zhang et al. (2023a); Wang et al. (2022) to incorporate the feedback from compilers to prompt the code generation ability of LLMs. Instead of directly feeding the code bug message to LLMs, we design the interactive chain of repair mechanism, which builds two agents to rethink and repair the code errors. The compiler serves as an INTERVENOR to avoid the Degeneration-of-Thought (DoT) problem during the interactive repair process.
3.2 Interactive Chain-of-Repair (CoR)
As shown in Figure 2, given the code generation tasks, INTERVENOR aims to mimic the human bug-repairing behavior by iteratively acquiring feedback from compilers, then generating bug reasons and solving plans in natural language, and finally fixing the program. Specifically, INTERVENOR employs two LLM based agents to play different roles in code repair (Sec. 3.2.1). Then we conduct an interactive code repair process using the agents (Sec. 3.2.2).
@@ 3.2.1 Agent Building
INTERVENOR involves the integration of two agents, Code Learner and Code Teacher, who work interactively to repair the generated codes. The role of the Code Learner follows the instructions to conduct code generation/repair, guided by the Code Teacher. The primary focus of the Code Teacher is to rethink and elucidate code errors for students. Additional details about agent construction can be found in Appendix A.5.
Code Learner. The Code Learner follows instructions and engages in two coding tasks, including initial code generation and code repair.
In the initial code generation phase, the Code Learner endeavors to generate the initial version of code according to the requirements of the given coding task. Subsequently, the agent’s role is modified for code repair. We trigger the code repair proficiency of LLMs using the instruction “You are a student assistant with excellent code repair capabilities”. Besides, Code Learner incorporates the Chain-of-Repair (CoR) as an instruction to guide the code repair.
Code Teacher. The Code Teacher conducts the CoR results during interactive code repair. It aims to generate the code error explanation and give bug-fixing planning for Code Learner.
Specifically, we use the prompt–“You are an experienced and insightful programming instructor”–to instruct LLMs to function as proficient code debuggers and serve to activate their bug-tracing and code-diagnosis abilities. Code Teacher integrates feedback from code compilers to produce extensive repair suggestions and guidance. This assists Code Learner in gaining a deeper understanding and effectively addressing errors within their code.
3.2.2 Interactive Code Repair Workflow
INTERVENOR conducts an interactive code-repair process, facilitating the collaboration among agents and the code compiler.
@@ 4 Experimental Methodology
In this section, we describe the datasets, evaluation metrics, baselines, and implementation details.
@@
Table 1: Data Statistics. #Tests represents the average number of test cases.
@@
Dataset. We evaluate the code generation and translation effectiveness on three datasets, including HumanEval, MBPP and HumanEval-X. Besides, we build a new benchmark CodeError to further test the code repair ability of LLMs. All data statistics are shown in Table 1.
@@
HumanEval Chen et al. (2021) serves as a benchmark for evaluating the functional correctness of synthesized programs generated from docstrings. It comprises 164 hand-written Python programming problems, which consist of function signatures, docstrings, bodies, and multiple unit tests. MBPP Austin et al. (2021) is a benchmark that includes 974 introductory-level Python programming problems. Each problem comprises a problem statement, a code solution, and three automated test cases, and the task IDs that range from 11 to 510 are used for evaluation. HumanEval-X Zheng et al. (2023) is used to assess a model’s multi-programming language generation and translation capability. It consists of 820 human-crafted data instances, covering C++, Java, JavaScript, and Go.
Then we build the CodeError benchmark to further evaluate the code repair effectiveness of INTERVENOR. The CodeError benchmark contains a total of 4,463 examples, evenly distributed across more than six different error types. It includes basic programming problems, data analysis problems, and programming competition problems. For each example, there are 9 test cases on average to evaluate the code’s correctness. More details of CodeError are shown in Appendix A.3.
@@
Evaluation Metrics. We use Pass@$k$ Chen et al. (2021) to evaluate the effectiveness of different models on both code generation task and code translation task, which is the same as the previous work Chen et al. (2021); Zheng et al. (2023); Li et al. (2023a); Chen et al. (2023a); Nijkamp et al. (2023).
@@
Baselines. We first compare INTERVENOR with several code-oriented large language pretrained models, such as Incoder Fried et al. (2023), CodeGen Nijkamp et al. (2023), CodeGeeX Zheng et al. (2023), CodeT5 Wang et al. (2023b), StarCoder Li et al. (2023b), WizardCoder Luo et al. (2023), and Llama based models Touvron et al. (2023); Roziere et al. (2023). These models are pretrained on large-scale code corpora, demonstrating strong code generation capabilities. Additionally, we also compare INTERVENOR with some closed-source and high-performance large language models, e.g. Claude Anthropic (2023), GPT-3.5 OpenAI (2022), and GPT-4 OpenAI (2023), which show strong emergent abilities, especially for the code generation tasks. In our experiments, GPT-3.5 is our main baseline model. Besides, we also compare Self-Debug Chen et al. (2023b) and the multi-agent collaborative method, Self-Collaboration Dong et al. (2023) to show the code repair ability of INTERVENOR.
@@
Implementation Details. In our experiments, we use GPT-3.5 as the foundation model to build different agents in our INTERVENOR model. We set the temperature to 0.2 and the maximum generation length to 512 tokens. The maximum number of interactive code repairs is set to 5. Additionally, we also use CodeLlama-7B/13B to implement the agents, Code Learner and Code Teacher, of our INTERVENOR model to explore the impact of using different LLMs.